*Problem 4-47A

Jessica Company manufactures hockey pucks and soccer balls. For both products, materials are added at the beginning of the production process and conversion costs are incurred evenly. Jessica uses the FIFO method to calculate equivalent units. Production and cost data for the month of August are as follows:
Production Data—Hockey pucks Units Percent Complete
Work in process units, August 1 400 70%
Units started into production 1,550
Work in process units, August 31 580 30%
Cost Data—Hockey pucks
Work in process, August 1 $1,040
Direct materials 1,550
Direct labour 1,130
Manufacturing overhead 892
Production Data—Soccer balls
Work in process units, August 1 270 90%
Units started into production 2,020
Work in process units, August 31 200 60%
Cost Data—Soccer balls
Work in process, August 1 $440
Direct materials 3,232
Direct labour 970
Manufacturing overhead 1,784
Calculate the following for both the hockey pucks and the soccer balls: (Round answers to the nearest whole dollar, e.g. 5,275. Round per unit costs to the 3 decimal places, e.g. 15.253.)

1. The equivalent units of production for materials and conversion costs.
Hockey pucks Soccer balls
Materials

Conversion Cost


2. The unit costs of production for materials and conversion costs.
Hockey pucks Soccer balls
Materials $

$

Conversion Cost $

$


3. The assignment of costs to units transferred out and to work in process at the end of the accounting period.
Hockey pucks Soccer balls
Beginning WIP $

$

Complete beginning WIP $

$

Started and completed $

$

Ending WIP $

$

Prepare a production cost report for the month of August for the hockey pucks only. (Round unit cost to 3 decimal places, e.g. 15.251 and other answers to to the nearest whole dollar, e.g. 5,275.)
JESSICA COMPANY
Production Cost Report—Hockey pucks
For the Month Ended August 31
Equivalent Units
Quantities Physical Units Materials Conversion Costs
Units to be accounted for
Work in process, August 1

Started into production

Total units

Units accounted for
Completed and transferred out
Work in process, August 1

Started and completed

Work in process, August 31

Total units

Costs
Unit costs
Costs in August $

$

Equivalent units

Unit costs $

$

Costs to be accounted for
Work in process, August 1 $

Added into production

Total costs $
Answer #1
Calculate the following for both the hockey pucks and the soccer balls: (Round answers to the nearest whole dollar, e.g. 5,275. Round per unit costs to the 3 decimal places, e.g. 15.253.)

1. The equivalent units of production for materials and conversion costs.
Hockey pucks Soccer balls
Materials

1550

2020

Conversion Cost

(400*30%+970+580*30%) = 1264

(270*10%+1820+200*60%) = 1967


2. The unit costs of production for materials and conversion costs.
Hockey pucks Soccer balls
Materials $1550/1550 = 1

$3232/2020 = 1.60

Conversion Cost $2022/1264 = 1.60

$2754/1967 = 1.40


3. The assignment of costs to units transferred out and to work in process at the end of the accounting period.
Hockey pucks Soccer balls
Beginning WIP $1040

$440

Complete beginning WIP $120

$43

Started and completed $970*2.6 = 2522

$5460

Ending WIP $858 $488

Prepare a production cost report for the month of August for the hockey pucks only. (Round unit cost to 3 decimal places, e.g. 15.251 and other answers to to the nearest whole dollar, e.g. 5,275.)
JESSICA COMPANY
Production Cost Report—Hockey pucks
For the Month Ended August 31
Equivalent Units
Quantities Physical Units Materials Conversion Costs
Units to be accounted for
Work in process, August 1

400

Started into production

1550

Total units

1950

Units accounted for
Completed and transferred out
Work in process, August 1

400

0

400*30% = 120

Started and completed

970

970

970

Work in process, August 31

580

580

580*30% = 174

Total units

1950

1550

1264

Costs
Unit costs
Costs in August $1550

$2022
Equivalent units

1550

1264

Unit costs $1

$1.60
Costs to be accounted for
Work in process, August 1 $1040

Added into production

3572

Total costs $4612